OpenStudy (anonymous):

The camera Melissa wanted for her birthday is on sale 28% off the usual price.The amount of the discont is $100.80.What was the original price of the camera?

OpenStudy (kropot72):

$100.80 is 28% of the usual price. Therefore 1% of the usual price is\[\frac{100.80}{28}\] and 100% of the usual (original) price is \[\frac{100.80}{28}\times 100=$\ (you\ can\ calculate)\ \]

OpenStudy (anonymous):

is the final $140

OpenStudy (anonymous):

answer*

OpenStudy (kropot72):

Did you do the calculation that I gave you?